OpenStudy (anonymous):

What is the center and radius of a circle with the given equation? x^2 +y^2 -4x + 10y =7

OpenStudy (anonymous):

Please HELP!!

OpenStudy (anonymous):

(x^2-4x)+(y^2+10y)=7 complete the squares and write in the form (x-h)^2+(y-k)^2=r^2 then (h,k) is the centre and r is the radius.